WATHCING THE PROTEIN MAMBO: FAST ENZYME DYNAMICS

\underline{CHRISTOPHER M. CHEATUM}, Department of Chemistry, University of Iowa, Iowa City, IA 52245.

The structural dynamics of enzymes at the femtosecond to picosecond time scale have been invoked to explain the results of temperature-dependent kinetic-isotope-effect measurements for a number of enzymatic reactions. We report studies of enzyme-ligand interaction dynamics at this time scale. We report studies of enzyme dynamics in complexes of formate dehydrogenase including transition-state-analog complexes and identify distinct and unexpected dynamics associated with the transition-state structure. Our results support a potential role for fast dynamics near the transition state and reveal differences in the nature of enzyme-ligand interaction dynamics in the ground state and in vicinity of the transition state of a reaction.